# Debias your Large Multi-Modal Model at Test-Time via Non-Contrastive Visual Attribute Steering

### v2

@@ 1 Introduction
Deep neural networks are well known to exhibit a myriad of societal biases learned from their training datasets [7, 45]. While this effect can be observed to some extent for most models, it is especially prevalent in models trained on web-scale data. Numerous prior works have observed and identified such biases present in modern Large Language Models (LLMs) [4, 8], while other recent work shows that such biases are even more prevalent in Large Multimodal Models (LMMs)  [^1] LMMs refer specifically to generative multimodal models. We refer to the general class of vision-language models as VLMs [6] such as LLaVA-1.5 [24], or Llama-3.2 [11] that integrate a vision backbone with a pretrained LLM. Given that LLMs are often pretrained on relatively uncurated web-scale data [30], any further integration with additional models will inherit the particular biases of the LLM. Without additional safety tuning, these pre-existing biases may be amplified further when an LLM is augmented with pretrained visual capabilities [17], which also come with a distinct set of implicit societal biases present in the visual pretraining data. Evaluating and mitigating potentially harmful behaviors induced by these societal biases is becoming increasingly important in order to safely deploy multimodal generative AI systems that utilize LMMs.
@@
Figure 1: Our proposed approaches for debiasing LMMs. Left: our PARzero approach solves for a steering vector with a single optimization step toward the target attributes, requiring no additional data. Middle: our PAR approach constructs a steering vector from a dataset of contrasting activations given by attribute-related Q&A. Right: LMMs with basic debiasing mechanisms like reprompting are not effective for debiasing on real world data.
@@
Specifically, we propose a novel method for identifying and ablating (removing) biased representations within LMMs. Recent work has shown that LLMs tend to encode high-level concepts as linear directions in feature space [35], making it possible to steer a specific feature’s contribution to the model’s output. For example, adding a “steering vector” can amplify a desired feature [47], whereas ablating that direction can remove the feature [3]. We introduce two methods for constructing steering vectors: (1) a dataset-based approach, where steering vectors are computed offline by comparing model activations from contrasting datasets—one that provokes the feature of interest and one that does not, and (2) a novel optimization-based approach, which constructs the steering vector at inference time using just the input image and the target attribute for removal.
@@ 2 Related Work
Templeton et al. [35] showed that interpretable features can be found by training a sparse autoencoder [19] to decompose the activations of an LLM into separable features. These features were found to be monosemantic – able to locally augment downstream outputs when manually introduced during inference. Turner et al. [37] showed that a steering vector for amplifying behavior in LLMs could be constructed with a single pair of contrasting prompts, and Arditi et al. [3] demonstrated that ablation of an abstract concept like refusal in LLMs can be performed via dataset-based steering.
@@
While several approaches have been proposed for mitigating social biases in VLMs [39, 5, 43, 31, 10, 33, 18], prior research on addressing such biases in LMMs is lacking. Sathe et al. [28] and Fraser and Kiritchenko [12] utilized synthetically generated images to analyze the presence of bias in LMMs, but do not address bias mitigation strategies. Hinck et al. [16] employ steering-based strategies for improving the linguistic fidelity of LMMs, but do not consider other kinds of bias. Howard et al. [17] also leveraged synthetic images from the SocialCounterfactuals dataset [18] to measure bias in LMMs but at a much larger scale, finding that LMMs possess more bias than the corresponding LLM from which they were trained. While representation-based steering for reducing societal biases has been demonstrated in LLMs such as Claude 3 [35], our work is the first to demonstrate successful inference-time representation steering for reducing bias in LMMs.
@@ 3 Methodology
Motivated by prior work on steering LLM behavior via intervening on representations, we propose an approach for debiasing LMMs by identifying and ablating a linear direction in the LMM’s activation space that corresponds to its propensity to mention a set of target protected attributes. We propose two methods for achieving protected attribute removal: PAR and PARzero. PAR leverages a synthetic dataset of biased and neutral inputs to induce the target representations. For low resource environments, or for increased flexibility, PARzero requires no additional data, and steers by taking a single step of gradient descent toward the target attribute. Both approaches compute a steering vector by taking the difference between internal representations when the model is generating text with and without protected attribute mentions.
@@ 3.2.1 Evaluation datasets
To evaluate how open-ended LMM generations vary in response to images depicting people of different social attributes, we utilize images from three datasets which possess social attribute annotations: SocialCounterfactuals [18], DA-COCO [9, 44], and FACET [15]. We pair each dataset with our open-ended prompts to evaluate our approach. The SocialCounterfactuals dataset contains 171k synthetic images depicting people of various races, genders, and physical characteristics such as age and body type. Images are grouped into counterfactual sets, in which each image is highly similar in its depiction of a common subject (e.g., a doctor) while varying only according to intersectional social attributes (e.g., race & gender). We use subsets of the SocialCounterfactuals dataset for constructing steering directions and evaluating models for bias related to race, gender, age, and body type. The Demographic Annotations on COCO (DA-COCO) dataset contains 28k images of people with annotations for skin tone and gender, which we use for evaluating bias related to perceived race and gender. Finally, the FACET dataset contains 32k diverse, high-resolution images annotated by experts, covering 13 person attributes and 52 person classes. For our evaluation, we utilize the perceived age, race, and gender attribute labels. See section (9) of the supplementary material for additional evaluation details.
@@ 4.1 Experimental details
We conduct experiments by generating responses from two LMMs with and without our steering methods: LLaVA 1.5 [23] and Llama-3.2-Vision-Instruct [11]. We generated a total of 1,820,904 responses across the five open-ended prompts, three image datasets, two LMMs, two steering methods and three baselines evaluated in our experiments. Our generation experiments were conducted on 512 Intel® Gaudi 2 AI accelerators in the Intel® Tiber AI Cloud.
@@
For PAR, we compute steering directions as described in section (3.1) with respect to contrasting activations from generations containing protected attribute mentions, and benign text generations. To obtain generations that mention protected attributes, we pair targeted prompts with 256 synthetic images. For benign text we sample 256 image-prompt pairs from the LLaVA-Instruct-80K dataset [24], excluding instances with protected attributes. For PARzero we only make use of the input image, as well as a list of attribute-related tokens to construct the steering vector. For both PAR and PARzero we follow prior work [37, 34, 25] by computing the steering direction with respect to a layer near the middle of the language model. Further details of the experimental design can be found in section (8) of the supplementary material.
@@ 4.2 Protected Attribute Evaluation
Table (1) presents our main results for evaluating how well our proposed steering methods reduce a LMM’s tendency to mention protected attributes. Across all three datasets, both PAR and PARzero lower both LLaVA and Llama’s propensity to generate target-attribute text compared to the unsteered baseline, as confirmed by both the LLM-as-judge and bigram count metrics. PAR and PARzero also improve over prompt-based baselines in nearly all conditions under the bigram count metric, as well as in all conditions with LLaVA under the LLM-as-judge metric. We observe that Llama-3.2 tends to respond well to the prompt-based baselines under the LLM-as-judge metric, especially for synthetic data e.g. SocialCounterfactualsWe hypothesize that the extensive chat tuning of Llama (relative to Llava 1.5) enables it to respond in a manner that the satisfies the GPT4o judge. We also believe that steering Llama-style LMMs may present additional challenges related to their per-layer vision-language fusion. Finally, we note that PARzero improves over the “Unsteered” baseline in nearly all conditions. PARzero does not require iterative prompting like the baselines, or an offline computation stage like PAR, making it the most efficient option for reducing protected attribute mentions in LMMs. We find that PARzero makes the greatest gains over the “Unsteered” baseline real data (FACET, DA-COCO) with Llama-3.2. This possibly signals that stronger models like Llama-3.2 PARzero is a
@@ 4.5 Fidelity of Steered Models to Image Content
Given the previous results, its an open question if general captioning capabilities are degraded by steering. To investigate this we employed the LLM-as-a-judge approach [46] to evaluate whether steering affects the accuracy of generated responses. We used GPT-4o to evaluate whether LLaVA’s text responses, with and without steering, match the corresponding image. GPT-4o was given the image and prompt: “Does the description match the image? Answer with Yes or No.” Manual analysis showed that GPT-4o responds “No” when the LMM hallucinates image details. The results in table (4) show no significant difference in accuracy between baseline and steered LLaVA models, indicating that steering does not degrade performance.
@@ 6 Discussion
Despite our best efforts to improve the fairness of generative AI models, we acknowledge that our choice of models, methodologies, and datasets may themselves contain latent biases which limit our ability to address this multi-faceted problem. The datasets employed in our evaluations as well as our automated evaluation methods which employ LLMs may themselves contain biases which could influence our results. Nevertheless, we feel that our manual validations and the scale of the study we conducted necessitated the use of these resources. We also acknowledge that our study was only conducted with the English language. Additional work is needed to explore the effectiveness of our steering methods for other languages. Despite these limitations, we hope that the findings of our study will raise awareness of the need for methods to debias LMMs and promote the development of more responsible and fair generative AI models.
@@
Finally, our single-gradient-step approach PARzero offers a promising framework for further development in bias mitigation and model alignment. This initial study suggests a strong foundation for rapid intervention in model behavior, which could be extended to other forms of model control. We also see an opportunity to explore techniques for steering different LMM architectures, given the different performance of LLaVA vs Llama. With the success observed in attribute suppression, further research may explore other types of steering or alignment tasks that could similarly benefit from a minimally invasive, gradient-based intervention. Next steps should aim to expand bias mitigation techniques to encompass a broader spectrum of attributes and assess the long-term impacts of steering interventions on model performance.
@@ 7 PARzero Loss Analysis
This results in a cumulative gradient heavily influenced by $-p_{j}$, yielding a relatively small gradient magnitude when $p_{j}$ values are spread close to uniform over $\mathcal{T}$. Consequently, KL-Divergence provides weak gradients in settings where uniform probability across the target tokens is desired, limiting its effectiveness.

### v3

@@ 1 Introduction
Deep neural networks are well known to exhibit a myriad of societal biases learned from their training datasets [7, 45]. While this effect can be observed to some extent for most models, it is especially prevalent in models trained on web-scale data. Numerous prior works have observed and identified such biases present in modern Large Language Models (LLMs) [4, 8], while other recent work shows that such biases are even more prevalent in Large Multimodal Models (LMMs)  [^2] LMMs refer specifically to generative multimodal models. We refer to the general class of vision-language models as VLMs [6] such as LLaVA-1.5 [24], or Llama-3.2 [11] that integrate a vision backbone with a pretrained LLM.
Given that LLMs are often pretrained on relatively uncurated web-scale data [30], any further integration with additional models will inherit the particular biases of the LLM. Without additional safety tuning, these pre-existing biases may be amplified further when an LLM is augmented with pretrained visual capabilities [17], which also come with a distinct set of implicit societal biases present in the visual pretraining data. Evaluating and mitigating potentially harmful behaviors induced by these societal biases is becoming increasingly important in order to safely deploy multimodal generative AI systems that utilize LMMs.
@@
Figure 1: Our approaches for debiasing LMMs. Left: our PARzero approach constructs a steering vector with a single optimization step toward the target attributes, requiring no additional data. Middle: our PAR approach constructs a steering vector from a dataset of activations given by attribute-related Q&A. Right: LMMs with basic debiasing mechanisms are ineffective on real world data.
@@
Specifically, we propose a novel method for identifying and ablating (removing) biased representations within LMMs. Recent work has shown that LLMs tend to encode high-level concepts as linear directions in feature space [35], making it possible to steer a specific feature’s contribution to the model’s output. For example, adding a “steering vector” can amplify a desired feature [47], whereas ablating that direction can remove the feature [3].
We introduce two methods for constructing steering vectors: (1) a dataset-based approach, where steering vectors are computed offline by comparing model activations from contrasting datasets—one that provokes the feature of interest and one that does not, and (2) a novel optimization-based approach, which constructs the steering vector at inference time using just the input image and the target attribute for removal.
@@ 2 Related Work
Templeton et al. [35] showed that interpretable features can be found by training a sparse autoencoder [19] to decompose the activations of an LLM into separable features. These features were found to be monosemantic – able to locally augment downstream outputs when manually introduced during inference. Turner et al. [37] showed that a steering vector for amplifying behavior in LLMs could be constructed with a single pair of contrasting prompts, and Arditi et al. [3] demonstrated that ablation of concepts like refusal can be performed via dataset-based steering.
@@
While several approaches have been proposed for mitigating social biases in VLMs [39, 5, 43, 31, 10, 33, 18], prior research on addressing such biases in LMMs is lacking. Sathe et al. [28] and Fraser and Kiritchenko [12] utilized synthetically generated images to analyze the presence of bias in LMMs, but do not address bias mitigation strategies. Hinck et al. [16] employ steering-based strategies for improving the linguistic fidelity of LMMs, but do not consider other kinds of bias. Howard et al. [17] also leveraged synthetic images from the SocialCounterfactuals dataset [18] to measure bias in LMMs but at a much larger scale, finding that LMMs possess more bias than the corresponding LLM from which they were trained.
While representation-based steering for reducing societal biases has been demonstrated in LLMs such as Claude 3 [35], our work is the first to demonstrate successful inference-time representation steering for reducing bias in LMMs.
@@ 3 Methodology
Motivated by prior work on steering LLM behavior via intervening on representations, we propose an approach for debiasing LMMs by identifying and ablating a linear direction in the LMM’s activation space that corresponds to its propensity to mention a set of target protected attributes. We propose two methods for achieving protected attribute removal: PAR and PARzero. PAR leverages a synthetic dataset of biased and neutral inputs to induce the target representations. For low resource environments, or for increased flexibility, PARzero requires no additional data, and steers by taking a single step of gradient descent toward the target attribute.
Both approaches compute a steering vector by taking the difference between internal representations when the model is generating text with and without protected attribute mentions.
@@ 3.2.1 Evaluation datasets
To evaluate how open-ended LMM generations vary in response to images depicting people of different social attributes, we utilize images from three datasets which possess social attribute annotations: SocialCounterfactuals [18], DA-COCO [9, 44], and FACET [15]. We pair each dataset with our open-ended prompts to evaluate our approach. The SocialCounterfactuals dataset contains 171k synthetic images depicting people of various races, genders, and physical characteristics such as age and body type. Images are grouped into counterfactual sets, in which each image is highly similar in its depiction of a common subject (e.g., a doctor) while varying only according to intersectional social attributes (e.g., race & gender).
We use subsets of the SocialCounterfactuals dataset for constructing steering directions and evaluating models for bias related to race, gender, age, and body type. The Demographic Annotations on COCO (DA-COCO) dataset contains 28k images of people with annotations for skin tone and gender, which we use for evaluating bias related to perceived race and gender. Finally, the FACET dataset contains 32k diverse, high-resolution images annotated by experts, covering 13 person attributes and 52 person classes. For our evaluation, we utilize the perceived age, race, and gender attribute labels. See section (9) of the supplementary material for additional evaluation details.
@@ 4.1 Experimental details
We conduct experiments by generating responses from two LMMs with and without our steering methods: LLaVA 1.5 [23] and Llama-3.2-Vision-Instruct [11].
We generated a total of 1,820,904 responses across the five open-ended prompts, three image datasets, two LMMs, two steering methods and three baselines evaluated in our experiments. Our generation experiments were conducted on 512 Intel® Gaudi 2 AI accelerators in the Intel® Tiber AI Cloud.
@@
For PAR, we compute steering directions as described in section (3.1) with respect to contrasting activations from generations containing protected attribute mentions, and benign text generations. To obtain generations that mention protected attributes, we pair targeted prompts with 256 synthetic images. For benign text we sample 256 image-prompt pairs from the LLaVA-Instruct-80K dataset [24], excluding instances with protected attributes. For PARzero we only make use of the input image, as well as a list of attribute-related tokens to construct the steering vector. For both PAR and PARzero we follow prior work [37, 34, 25] by computing the steering direction with respect to a layer near the middle of the language model.
@@ 4.2 Protected Attribute Evaluation
Table (1) presents our main results for evaluating how well our proposed steering methods reduce a LMM’s tendency to mention protected attributes. Across all three datasets, both PAR and PARzero lower both LLaVA and Llama’s propensity to generate target-attribute text compared to the unsteered baseline, as confirmed by both the LLM-as-judge and bigram count metrics. PAR and PARzero also improve over prompt-based baselines in nearly all conditions under the bigram count metric, as well as in all conditions with LLaVA under the LLM-as-judge metric. We observe that Llama-3.2 tends to respond well to the prompt-based baselines under the LLM-as-judge metric, especially for synthetic data e.g. SocialCounterfactualsWe hypothesize that the extensive chat tuning of Llama (relative to Llava 1.5) enables it to respond in a manner that the satisfies the GPT4o judge. We also believe that steering Llama-style LMMs may present additional challenges related to their per-layer vision-language fusion. Finally, we note that PARzero improves over the “Unsteered” baseline in nearly all conditions. PARzero does not require iterative prompting like the baselines, or an offline computation stage like PAR, making it the most efficient option for reducing protected attribute mentions in LMMs.
@@ 4.5 Fidelity of Steered Models to Image Content
Given the previous results, its an open question if general captioning capabilities are degraded by steering. To investigate this we employed the LLM-as-a-judge approach [46] to evaluate whether steering affects the accuracy of generated responses. We used GPT-4o to evaluate whether LLaVA’s text responses, with and without steering, match the corresponding image.
Manual analysis showed that GPT-4o responds “No” when the LMM hallucinates image details. The results in table (4) show no significant difference in accuracy between baseline and steered LLaVA models.
@@ 6 Discussion
Despite our best efforts to improve the fairness of generative AI models, we acknowledge that our choice of models, methodologies, and datasets may themselves contain latent biases which limit our ability to address this multi-faceted problem. The datasets employed in our evaluations as well as our automated evaluation methods which employ LLMs may themselves contain biases which could influence our results. Nevertheless, we feel that our manual validations and the scale of the study we conducted necessitated the use of these resources. We also acknowledge that our study was only conducted with the English language. Additional work is needed to explore the effectiveness of our steering methods for other languages. Despite these limitations, we hope that the findings of our study will raise awareness of the need for methods to debias LMMs and promote the development of more responsible and fair generative AI.
@@
This initial study suggests a strong foundation for rapid intervention in model behavior, which could be extended to other forms of model control. We also see an opportunity to explore techniques for steering different LMM architectures, given the different performance of LLaVA vs Llama. With the success observed in attribute suppression, further research may explore other types of steering or alignment tasks that could similarly benefit from a minimally invasive, gradient-based intervention.
@@ 7 PARzero Loss Analysis
This results in a cumulative gradient heavily influenced by $p_{j}$, yielding a relatively small gradient magnitude when $p_{j}$ values are spread close to uniform over $\mathcal{T}$. Consequently, KL-Divergence provides weak gradients in settings where uniform probability across the target tokens is desired, limiting its effectiveness.